# Euler-type approximation for the invariant measure: An abstract framework

### v1

@@ 2.1. Coupling
We suppose that
$\overline{\omega}(\pi):=\limsup_{n\rightarrow\infty}\frac{\gamma_{n}-\gamma_{n+1}}{\gamma_{n+1}^{2}}\in(0,\infty).$
@@
We need the following technical lemma.
@@ A. One may find $C_{b,\varepsilon}$ such that for all $n\in\mathbb{N}^{*}$
$\sigma_{b,\varepsilon}(n)$ $\leq$ $C_{b,\varepsilon}\gamma_{n}^{\varepsilon}\quad\text{if}\quad\frac{b}{\varpi(\pi)}>\varepsilon$ (2.8)
$\leq$ $C_{b,\lambda}\gamma_{n}^{\lambda}\quad\text{if}\quad\lambda<\frac{b}{\varpi(\pi)}\leq\varepsilon.$ (2.9)
@@ B. For the particular case $\gamma_{n}=\frac{1}{n+h^{-1}}$, one has $\varpi(\pi)=1$ and
Lemma 2.6.
Let $\Theta^{1}$ and $\Theta^{2}$ be families which are $(p,b_{\ast},\varepsilon)$-coupled for some $b_{\ast},\varepsilon>0$. We note $C_{\ast}>0$, $h\in(0,1)$ some constants such that (2.5) holds for $\Theta_{s,t}^{1}$ and $\Theta_{s,t}^{2}$, $s\leq t\leq s+h$. We also assume that $\Theta^{1}$ is $p$-bounded and $h\in(0,1)$ is sufficiently small to have $M_{p}(\Theta^{1},\mu,h)<\infty$ and, with $\gamma_{n}=\frac{1}{n+h^{-1}}$,
$2^{p}C_{\ast}\sigma_{b_{\ast},\varepsilon}(k)\leq\frac{1}{2},\quad k\leq n.$ (2.11)
Then, for every $n\in{\mathbb{N}}$, we have
Proof.
Step 1. We denote $a_{k}=$ $W_{p}^{p}(\Theta_{t_{0},t_{k}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{k}}^{2,\pi}(\mu^{2}))$ and we write, for $n\geq 1$,
We define
$M_{n}=\max_{k\leq n}\left\|\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1})\right\|_{p}^{p}\leq M_{p}(\Theta^{1},\mu^{1},h).$
Using the coupling property (2.5) first and (2.2) then, we obtain
$e^{-b_{\ast}(t_{n}-t_{k})}a_{k}-e^{-b_{\ast}(t_{n}-t_{k-1})}a_{k-1}$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{k})}C_{\ast}\left(1+2^{p}\left\|\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1})\right\|_{p}^{p}+2^{p}W_{p}^{p}(\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{k-1}}^{2,\pi}(\mu^{2}))\right)\gamma_{k}^{1+\varepsilon}$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{k})}C_{\ast}\left(1+2^{p}M_{n}+2^{p}a_{k-1}\right)\gamma_{k}^{1+\varepsilon}.$
Taking the sum, we get
$a_{n}$ $\leq$ $e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+C_{\ast}\left(1+2^{p}M_{n}\right)\sum_{k=1}^{n}e^{-b_{\ast}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}+C_{\ast}2^{p}\sum_{k=1}^{n}e^{-b_{\ast}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}a_{k-1}$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+C_{\ast}\left(1+2^{p}M_{n}+2^{p}\max_{k\leq n-1}a_{k}\right)\sigma_{b_{\ast},\varepsilon}(n)$
$\leq$ $a_{0}+\frac{1}{2}\left(2^{-p}+M_{n}\right)+\frac{1}{2}\times\max_{k\leq n-1}a_{k}$
the last inequality being a consequence of our hypothesis (2.11). This gives
$\max_{i\leq n}a_{i}\leq a_{0}+\frac{1}{2}\left(1+M_{n}\right)+\frac{1}{2}\times\max_{k\leq n-1}a_{k}$
and finally
$\max_{i\leq n}a_{i}\leq 2a_{0}+M_{n}+1$ (2.14)
Step 2.We come back to (2.1) and we use (2.2) and (2.14) in order to get
$e^{-b_{\ast}(t_{n}-t_{k})}a_{k}-e^{-b_{\ast}(t_{n}-t_{k-1})}a_{k-1}$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{k})}C_{\ast}\left(1+2^{p}M_{n}+2^{p}a_{k-1}\right)\gamma_{k}^{1+\varepsilon}$
$\leq$ $e^{-b_{\ast}(t_{n}-t_{k})}2^{p+1}C_{\ast}\left(1+M_{n}+a_{0}\right)\gamma_{k}^{1+\varepsilon}.$
Taking the sum and using Lemma 2.5, we conclude with
$a_{n}$ $\leq$ $e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+2^{p+1}C_{*}(1+M_{n}+a_{0})\sum_{k=1}^{n}e^{-b_{\ast}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}$
$=$ $e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+2^{p+1}C_{*}(1+M_{n}+a_{0})\sigma_{b_{\ast},\varepsilon}(n).$
Now, we observe that
$\max_{k\leq n}\left\|\Theta_{t_{0},t_{k-1}}^{2,\pi}(\mu)\right\|_{p}^{p}\leq 2^{p-1}M_{n}+2^{p-1}\max_{k\leq n}a_{k},$
which gives, by (2.14), $M_{p}(\Theta^{2},\mu,h)\leq 2^{p}M_{p}(\Theta^{1},\mu,h)+2^{p}a_{0}+2^{p-1}<\infty$ for every $\mu\in\mathcal{P}_{p}(B)$. ∎
@@
We suppose that the hypotheses of the previous lemma are in force and we consider the grid $t_{n}=t_{0}+\sum_{i=1}^{n}\frac{1}{i+h^{-1}}.$ Then, for $b_{*}\neq\varepsilon$,
$W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq$ $An^{-b_{\ast}\wedge\varepsilon},\$(2.15)
with $A=\left((2^{p+1}C_{*}C_{b_{\ast},\varepsilon}+1)W_{p}^{p}(\mu^{1},\mu^{2})+2^{p+1}C_{*}C_{b_{\ast},\varepsilon}(1+M_{p}(\Theta^{1},\mu^{1},h))\right)$. In the case $b=\varepsilon$, one has to multiply this upper bound by $1+\ln(1+n)$. [^2] In the same way, we have $W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq A(1+(t_{n}-t_{0}))e^{-b_{\ast}\wedge\varepsilon(t_{n}-t_{0})}.$
@@ 2.2. Invariant measure
where $A=\left((2^{p+1}C_{*}\tilde{C}_{b_{\ast},\varepsilon}+1)W_{p}^{p}(\mu^{1},\mu^{2})+2^{p+1}C_{*}\tilde{C}_{b_{\ast},\varepsilon}(1+M_{p}(\theta,\mu^{1}))\right)$ and $\tilde{C}_{b_{\ast},\varepsilon}=C_{b_{\ast},\varepsilon}+1$.
@@
Let $s<t$ and $\gamma_{k}=\frac{1}{k+h^{-1}},t_{k}=t_{0}+\gamma_{1}..+\gamma_{k}.$ Let $n\in{\mathbb{N}}$ be such that $t_{n}\leq t<t_{n+1}$. Then, we consider the grid $t_{k}^{\prime}=t_{k}$ for $k\in\{0,\dots,n\}$ and $t_{n+1}^{\prime}=t\leq t_{n+1}$. This means that $\gamma_{k}^{\prime}=\gamma_{k}$ for k=1,,nk=1,...,n and $\gamma_{n+1}^{\prime}<\gamma_{n+1}.$ Recall that, by (2.10), $\sigma_{b_{*},\varepsilon}(n)\leq C_{b_{*},\varepsilon}e^{-b_{*}\wedge\varepsilon(t-s)}(1+(t-s)^{1_{b_{*}=\varepsilon}})$ and by (A.2) $\gamma_{n}\leq h^{-1}e^{-(t_{n}-s).}.$ Then, if $\sigma_{b_{*},\varepsilon}^{\prime}(n)$ is associated to $\gamma_{i}^{\prime},i=0,1,$..., we have
@@
Consider $\theta_{s,t}:{\mathcal{P}}_{p}(B)\rightarrow{\mathcal{P}}_{p}(B)$ a continuous time-homogeneous flow, which is $p$-bounded and is $(p,b_{\ast},\varepsilon)$ self-coupled for some $h\in(0,1)$, $b_{\ast},\varepsilon>0.$ Then, there exists a unique probability measure $\nu\in\mathcal{P}_{p}(B)$ such that for every $s\leq t$ one has $\nu=\theta_{s,t}(\nu).$ It satisfies $\|\nu\|_{p}^{p}=M_{p}(\theta,\nu)$ and we have for every $s\leq t$ and every $\mu\in\mathcal{P}_{p}(B)$,
$W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq Ae^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s)^{\mathbf{1}_{b_{\ast}=\varepsilon}})$ (2.18)
with $A$ being as given by Corollary 2.9.
@@
$W_{p}^{p}(\theta_{s,t}(\mu),\theta_{s,t+\delta}(\mu))=W_{p}^{p}(\theta_{s,t}(\mu),\theta_{s,t}(\theta_{t,t+\delta}(\mu)))\leq Ae^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s))^{\mathbf{1}_{b_{\ast}=\varepsilon}}.$ (2.19)
Therefore, if $t_{n}\rightarrow\infty$ , we get that $\theta_{s,t_{n}}(\mu)$, $n\in{\mathbb{N}}$ is a Cauchy sequence. Since $\mathcal{P}_{p}(B)$ is complete with respect to $W_{p}$, we define its limit in the Wasserstein metric as $\nu:=\lim_{n\rightarrow\infty}\theta_{s,t_{n}}(\mu)\in\mathcal{P}_{p}(B)$. Coming back to 2.19 and passing to the limit we get
@@
Let $\theta$ be a time homogeneous continuous flow on ${\mathcal{P}}_{p}(B)$ which is $p$-bounded and is $(p,b_{\ast},\varepsilon)$ self-coupled for some $h\in(0,1)$, $b_{\ast},\varepsilon>0$ and $\nu$ the invariant measure obtained in Proposition 2.11. Let $\Theta_{s,t}:{\mathcal{P}}_{p}(B)\rightarrow{\mathcal{P}}_{p}(B)$, $s\leq t$, be a family of applications which is $(p,b_{\ast},\varepsilon)$ coupled with $\theta_{s,t}$. Let $h\in(0,1)$ be such that both (2.11) and the coupling property (2.5) between $\theta$ and $\Theta$ hold. Let $\gamma_{n}=\frac{1}{n+h^{-1}}$ and $t_{n}=s+\gamma_{1}+...+\gamma_{n}$, $n\in{\mathbb{N}}$. Then, we have for every $\mu\in\mathcal{P}_{p}(B)$
@@
with $A=2^{2p}C_{*}(1+M_{p}(\theta,\mu))(2C_{b_{\ast},\varepsilon}+1)$.
@@
$W_{p}^{p}(\Theta_{s,t_{n}}^{\pi}(\mu),\theta_{s,t_{n}}(\mu))\leq A^{\prime}n^{-b_{\ast}\wedge\varepsilon},$
with $A^{\prime}=2^{p+1}C_{*}C_{b_{\ast},\varepsilon}(1+M_{p}(\theta,\mu))$. By (2.18), we also have
$W_{p}^{p}(\theta_{s,t_{n}}(\mu),\nu)\leq A^{\prime\prime}e^{-\varepsilon\wedge b_{\ast}(t_{n}-s)}\leq A^{\prime\prime}n^{-\varepsilon\wedge b_{\ast}}$
with $A^{\prime\prime}=2^{p+1}C_{*}(C_{b_{\ast},\varepsilon}+1)(1+M_{p}(\theta,\mu))$, since $t_{n}-s\leq\ln(1+n)$. Using the triangular inequality and $(a+b)^{p}\leq 2^{p-1}(a^{p}+b^{p})$, we get the claim. ∎

### v2

@@ 2.1. Coupling on a time grid
Let us observe that $e^{-b(t_{n}-t_{i})}\gamma_{i}^{1+\varepsilon}\leq\gamma_{1}^{\varepsilon}\int_{t_{i-1}}^{t_{i}}e^{-b(t_{n}-s)}e^{b\gamma_{1}}ds$ because $\gamma$ is decreasing, and therefore
$\sigma_{b,\varepsilon}(n)\leq\gamma_{1}^{\varepsilon}\frac{e^{b\gamma_{1}}}{b}.$ (2.10)
Lemma 2.6.
Let $\Theta^{1}$ and $\Theta^{2}$ be families which are $(p,b_{\ast},\varepsilon)$-coupled for some $b_{\ast},\varepsilon>0$. We note $C_{\ast}>0$, $h\in(0,1)$ some constants such that (2.4) holds for $\Theta_{s,t}^{1}$ and $\Theta_{s,t}^{2}$, $s\leq t\leq s+h$. We also assume that $\Theta^{1}$ is $p$-bounded and $h\in(0,1)$ is sufficiently small to have $M_{p}(\Theta^{1},\mu,h)<\infty$ and
$2^{p}C_{\ast}h^{\varepsilon}\frac{e^{bh}}{b}\leq\frac{1}{2}.$ (2.11)
Then, for any positive decreasing sequence $(\gamma_{n})_{n\geq 1}$ such that $\gamma_{1}<h$ and $(t_{n})_{n\geq 0}$ be defined as in (2.8), we have for every $n\in{\mathbb{N}}$
Proof.
We denote $a_{k}=$ $W_{p}^{p}(\Theta_{t_{0},t_{k}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{k}}^{2,\pi}(\mu^{2}))$ and first prove an upper bound for this sequence. We write, for $n\geq 1$,
We first observe
$\forall n,\ \max_{k\leq n}\left\|\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1})\right\|_{p}^{p}\leq M_{p}(\Theta^{1},\mu^{1},h)=:{\bf M}.$
Using the coupling property (2.4) first and (2.2) then, we obtain
$e^{-b_{\ast}(t_{n}-t_{k})}a_{k}-e^{-b_{\ast}(t_{n}-t_{k-1})}a_{k-1}$ (2.13)
$\leq C_{\ast}e^{-b_{\ast}(t_{n}-t_{k})}\!\!\left(\!1+\!2^{p}\left\|\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1})\right\|_{p}^{p}\!+\!2^{p}W_{p}^{p}(\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{k-1}}^{2,\pi}(\mu^{2}))\!\right)\!\gamma_{k}^{1+\varepsilon}$
$\leq C_{\ast}e^{-b_{\ast}(t_{n}-t_{k})}\left(1+2^{p}{\bf M}+2^{p}a_{k-1}\right)\gamma_{k}^{1+\varepsilon}.$
Taking the sum, we get
$a_{n}$ $\leq e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+C_{\ast}\left(1+2^{p}{\bf M}\right)\sum_{k=1}^{n}e^{-b_{\ast}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}$
$\quad+C_{\ast}2^{p}\sum_{k=1}^{n}e^{-b_{\ast}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}a_{k-1}$
$\leq e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+C_{\ast}\left(1+2^{p}{\bf M}+2^{p}\max_{k\leq n-1}a_{k}\right)\sigma_{b_{\ast},\varepsilon}(n)$
$\leq a_{0}+\frac{1}{2}\left(2^{-p}+{\bf M}+\max_{k\leq n-1}a_{k}\right),$
the last inequality being a consequence of our hypothesis (2.11) and the upper bound (2.10). Taking the maximum, this gives
$\max_{i\leq n}a_{i}\leq a_{0}+\frac{1}{2}\left(1+{\bf M}\right)+\frac{1}{2}\times\max_{k\leq n-1}a_{k}$
and finally
$\max_{i\leq n}a_{i}\leq 2a_{0}+{\bf M}+1.$ (2.14)
We now come back to (2.13) and we use (2.2) and (2.14) in order to get
$e^{-b_{\ast}(t_{n}-t_{k})}a_{k}-e^{-b_{\ast}(t_{n}-t_{k-1})}a_{k-1}$ $\leq e^{-b_{\ast}(t_{n}-t_{k})}C_{\ast}\left(1+2^{p}{\bf M}+2^{p}a_{k-1}\right)\gamma_{k}^{1+\varepsilon}$
$\leq e^{-b_{\ast}(t_{n}-t_{k})}2^{p+1}C_{\ast}\left(1+{\bf M}+a_{0}\right)\gamma_{k}^{1+\varepsilon}.$
Taking the sum and using Lemma 2.7, we conclude with
$a_{n}$ $\leq e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+2^{p+1}C_{*}(1+{\bf M}+a_{0})\sum_{k=1}^{n}e^{-b_{\ast}(t_{n}-t_{k})}\gamma_{k}^{1+\varepsilon}$
$=e^{-b_{\ast}(t_{n}-t_{0})}a_{0}+2^{p+1}C_{*}(1+{\bf M}+a_{0})\sigma_{b_{\ast},\varepsilon}(n).$
Now, we observe that
$\max_{k\leq n}\left\|\Theta_{t_{0},t_{k-1}}^{2,\pi}(\mu)\right\|_{p}^{p}\leq 2^{p-1}\max_{k\leq n}\left\|\Theta_{t_{0},t_{k-1}}^{1,\pi}(\mu)\right\|_{p}^{p}+2^{p-1}\max_{k\leq n}a_{k}.$
Since the sequence $(t_{i})_{i\geq 0}$ is arbitrary among those such that $t_{1}-t_{0}<h$ and $(t_{i+1}-t_{i})_{i\geq 0}$ is decreasing, this gives by (2.14), $M_{p}(\Theta^{2},\mu,h)\leq 2^{p}M_{p}(\Theta^{1},\mu,h)+2^{p}a_{0}+2^{p-1}<\infty$ for every $\mu\in\mathcal{P}_{p}(B)$. ∎
To go further, we will make the additional assumption on $(\gamma_{n})$
$\varpi(\pi):=\limsup_{n\rightarrow\infty}\frac{\gamma_{n}-\gamma_{n+1}}{\gamma_{n+1}^{2}}\in[0,\infty),$ (2.15)
that is needed for the following technical lemma.
@@ A. One may find $C_{b,\varepsilon}$ such that for all $n\in\mathbb{N}^{*}$
$\sigma_{b,\varepsilon}(n)\leq\begin{cases}C_{b,\varepsilon}\gamma_{n}^{\varepsilon}\text{ if }b>\varepsilon\varpi(\pi),\\ C_{b,\lambda}\gamma_{n}^{\lambda}\text{ if }\varpi(\pi)>0\text{ and }\lambda<\frac{b}{\varpi(\pi)}\leq\varepsilon.\end{cases}$ (2.16)
@@ B. For the particular case $\gamma_{n}=\frac{1}{n+h^{-1}}$, one has $\varpi(\pi)=1$ and
Remark 2.8.
Let $\beta\in(0,1)$ and $\gamma_{n}=\frac{1}{n^{\beta}+h^{-1}}$. This sequence fulfills the required assumptions with $\varpi(\pi)=0$. Then, we have $t_{n}\sim\frac{n^{1-\beta}}{1-\beta}$ as $n\to\infty$. By Lemma 2.7, we have $\sigma_{b,\varepsilon}(n)\leq C_{b,\varepsilon}n^{-\beta\varepsilon}$. Besides, for $s<t$, $t_{n}=s+\sum_{i=1}^{n}\gamma_{i}$ and $n\in{\mathbb{N}}$ such that $t_{n}\leq t<t_{n+1}$, we have $n\sim\left((1-\beta)(t-s)\right)^{\frac{1}{1-\beta}}$, so that $\sigma_{b,\varepsilon}(n)\leq C(t-s)^{\frac{-\beta\varepsilon}{1-\beta}}$. As $t-s\to\infty$, this convergence is less sharp than (2.17). For this, and also to avoid multiple cases, we will only work with $\gamma_{n}=\frac{1}{n+h^{-1}}$ in the paper.
@@
We suppose that the hypotheses of Lemma 2.6 are in force and we consider the grid $t_{n}=t_{0}+\sum_{i=1}^{n}\frac{1}{i+h^{-1}}.$ Then, for $b_{*}\neq\varepsilon$, $\mu^{1},\mu^{2}\in\mathcal{P}_{p}(B)$,
$W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq$ $An^{-b_{\ast}\wedge\varepsilon},\$
with $A=\left((2^{p+1}C_{*}C_{b_{\ast},\varepsilon}+1)W_{p}^{p}(\mu^{1},\mu^{2})+2^{p+1}C_{*}C_{b_{\ast},\varepsilon}(1+M_{p}(\Theta^{1},\mu^{1},h))\right)$. In the case $b_{*}=\varepsilon$, one has to multiply this upper bound by $1+\ln(1+n)$. [^2] In the same way, we have $W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq A(1+(t_{n}-t_{0}))e^{-b_{\ast}\wedge\varepsilon(t_{n}-t_{0})}.$
@@
Remark 2.10.
In Lemma 2.6, the condition $\gamma_{1}<h$ is required to have $\gamma_{n}\in(0,h)$ for all $n$, which allows to use (2.4) for all time steps. This is why we then take $\gamma_{n}=\frac{1}{n+h^{-1}}$ in Corollary 2.9.
If the condition $\gamma_{1}<h$ is not satisfied, we can still define $n(h)=\inf\{n\geq 1:\gamma_{n}<h\}$ apply Lemma (2.6) (and Corollary 2.9) for $n\geq n(h)$ on the grid $\{t_{n(h)-1},\dots,t_{n}\}$. For example, let $\gamma_{n}=\frac{1}{a+n}$ for some $a\geq 0$ and $t_{n}=t_{0}+\sum_{i=1}^{n}\frac{1}{a+i}$. We have by Corollary 2.9
$W_{p}^{p}(\Theta_{t_{0},t_{n}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n}}^{2,\pi}(\mu^{2}))\leq A(n-n(h)+1)^{-b_{\ast}\wedge\varepsilon}\leq A^{\prime}n^{-b_{\ast}\wedge\varepsilon},$ (2.18)
with $A=(2^{p+1}C_{*}C_{b_{\ast},\varepsilon}+1)W_{p}^{p}(\Theta_{t_{0},t_{n(h)-1}}^{1,\pi}(\mu^{1}),\Theta_{t_{0},t_{n(h)-1}}^{2,\pi}(\mu^{2}))+2^{p+1}C_{*}C_{b_{\ast},\varepsilon}(1+M_{p}(\Theta^{1},\Theta_{t_{0},t_{n(h)-1}}^{1,\pi}(\mu^{1}),h))$ and $A^{\prime}=An(h)^{b_{\ast}\wedge\varepsilon}$.
Besides, if $\Theta^{1}$ and $\Theta^{2}$ satisfy the $p$-Foster Lyapunov criterion (3.2), we have $A^{\prime}\leq\tilde{A}^{\prime}\Gamma_{p}(\mu^{1},\mu^{2})$ for a constant $\tilde{A}^{\prime}\geq 0$ that does not depend on $(\mu^{1},\mu^{2})$.
@@ 2.2. Invariant measure
where $A=\left((2^{p+1}C_{*}\tilde{C}_{b_{\ast},\varepsilon}+1)W_{p}^{p}(\mu^{1},\mu^{2})+2^{p+1}C_{*}\tilde{C}_{b_{\ast},\varepsilon}(1+M_{p}(\theta,\mu^{1}))\right)$ and $\tilde{C}_{b_{\ast},\varepsilon}=C_{b_{\ast},\varepsilon}+1$. Besides, $A\leq C(\Gamma_{p}(\mu^{1},\mu^{2})+M_{p}(\theta,\mu^{1}))$ with $C$ that does not depend on $\mu^{1}$ and $\mu^{2}$.
Let us stress that this inequality is interesting only when tst-s is large as it gives an exponential rate of convergence, but is not sharp when $t\to s$.
@@
Let $s<t$ and $\gamma_{k}=\frac{1}{k+h^{-1}},t_{k}=t_{0}+\gamma_{1}+\dots+\gamma_{k}.$ Let $n\in{\mathbb{N}}$ be such that $t_{n}\leq t<t_{n+1}$. Then, we consider the grid $t_{k}^{\prime}=t_{k}$ for $k\in\{0,\dots,n\}$ and $t_{n+1}^{\prime}=t\leq t_{n+1}$. This means that $\gamma_{k}^{\prime}=\gamma_{k}$ for $k=1,\dots,n$ and $\gamma_{n+1}^{\prime}<\gamma_{n+1}.$ Recall that, by (2.17), $\sigma_{b_{*},\varepsilon}(n)\leq C_{b_{*},\varepsilon}e^{-b_{*}\wedge\varepsilon(t-s)}(1+(t-s)^{1_{b_{*}=\varepsilon}})$ and by (A.2) $\gamma_{n}\leq he^{-(t_{n}-s).}$. Then, with $\sigma_{b_{*},\varepsilon}^{\prime}(n)$ associated to $(\gamma_{i}^{\prime})_{i\geq 1}$, we have
@@
Consider $\theta_{s,t}:{\mathcal{P}}_{p}(B)\rightarrow{\mathcal{P}}_{p}(B)$ a continuous time-homogeneous flow, which is $p$-bounded and is $(p,b_{\ast},\varepsilon)$ self-coupled for some $h\in(0,1)$, $b_{\ast},\varepsilon>0.$ Then, there exists a unique probability measure $\nu\in\mathcal{P}_{p}(B)$ such that for every $s\leq t$ one has $\nu=\theta_{s,t}(\nu).$ It satisfies $\|\nu\|_{p}^{p}=M_{p}(\theta,\nu)$ and there exists $C\in\mathbb{R}_{+}$ such that for every $s\leq t$ and every $\mu\in\mathcal{P}_{p}(B)$,
$W_{p}^{p}(\theta_{s,t}(\mu),\nu)\leq C(1+M_{p}(\theta,\mu))e^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s)^{\mathbf{1}_{b_{\ast}=\varepsilon}}).$ (2.21)
@@
$W_{p}^{p}(\theta_{s,t}(\mu),\theta_{s,t+\delta}(\mu))$ $=W_{p}^{p}(\theta_{s,t}(\mu),\theta_{s,t}(\theta_{t,t+\delta}(\mu)))$
$\leq Ae^{-\varepsilon\wedge b_{\ast}(t-s)}(1+(t-s))^{\mathbf{1}_{b_{\ast}=\varepsilon}},$ (2.22)
with $A\leq C(1+M_{p}(\theta,\mu))$ and $C\in{\mathbb{R}}_{+}$ does not depend on $\mu$ (note that $\|\mu\|_{p}^{p}\leq M_{p}(\theta,\mu)$ since $\theta$ is a continuous flow.). Therefore, if $t_{n}\rightarrow\infty$ , we get that $\theta_{s,t_{n}}(\mu)$, $n\in{\mathbb{N}}$ is a Cauchy sequence. Since $\mathcal{P}_{p}(B)$ is complete with respect to $W_{p}$, we define its limit in the Wasserstein metric as $\nu:=\lim_{n\rightarrow\infty}\theta_{s,t_{n}}(\mu)\in\mathcal{P}_{p}(B)$. Coming back to (2.22) and passing to the limit we get
@@
Let $\theta$ be a time homogeneous continuous flow on ${\mathcal{P}}_{p}(B)$ which is $p$-bounded and is $(p,b_{\ast},\varepsilon)$ self-coupled for some $h\in(0,1)$, $b_{\ast},\varepsilon>0$ and $\nu$ the invariant measure obtained in Proposition 2.14. Let $\Theta_{s,t}:{\mathcal{P}}_{p}(B)\rightarrow{\mathcal{P}}_{p}(B)$, $s\leq t$, be a family of applications which is $(p,b_{\ast},\varepsilon)$ coupled with $\theta_{s,t}$. Let $h\in(0,1)$ be such that both (2.11) and the coupling property (2.4) between $\theta$ and $\Theta$ hold. Let $a\geq 0$, $\gamma_{n}=\frac{1}{a+n}$ and $t_{n}=s+\gamma_{1}+...+\gamma_{n}$, $n\in{\mathbb{N}}$. Then, for every $\mu\in\mathcal{P}_{p}(B)$, there exists $A\in{\mathbb{R}}_{+}$ such that
@@
where $A$ does not depend on $\mu$.
@@
$W_{p}^{p}(\Theta_{s,t_{n}}^{\pi}(\mu),\theta_{s,t_{n}}(\mu))\leq C(1+M_{p}(\theta,\mu)+\|\Theta^{\pi}_{s,t_{n(h)}}(\mu)\|_{p}^{p})n^{-b_{\ast}\wedge\varepsilon},$ (2.24)
noting that $M_{p}(\theta,\theta_{s,t}(\mu)))\leq M_{p}(\theta,\mu)$ since $\theta$ is a time homogeneous flow. By (2.21), we also have
$W_{p}^{p}(\theta_{s,t_{n}}(\mu),\nu)\leq C(1+M_{p}(\theta,\mu))e^{-\varepsilon\wedge b_{\ast}(t_{n}-s)}\leq C(1+M_{p}(\theta,\mu))n^{-\varepsilon\wedge b_{\ast}}.$
Using the triangular inequality and $(a+b)^{p}\leq 2^{p-1}(a^{p}+b^{p})$, we get the claim. ∎
Remark 2.16.
If we assume in addition that $\theta$ and $\Theta$ satisfy the p-Foster Lyapunov criterion (3.2), we see from the proof above that the constant $A$ in (2.23) satisfies $A\leq C(1+\|\mu\|_{p}^{p})$, with $C$ that does not depend on $\mu$, so that
$\forall\mu\in\mathcal{P}_{p}({\mathbb{R}}^{d}),\ W_{p}^{p}(\Theta_{s,t_{n}}^{\pi}(\mu),\nu)\leq C(1+\|\mu\|_{p}^{p})n^{-b_{\ast}\wedge\varepsilon}(1+\ln(n+1)^{\mathbf{1}_{b_{\ast}=\varepsilon}})$ (2.25)
Alternatively, if (only) $\theta$ satisfies the $p$-Foster Lyapunov criterion (3.2) and the parameter $a$ in Theorem 3.7 is such that $a\geq h^{-1}$, then $n(h)=0$ and $\Theta_{s,t_{n(h)}}^{\pi}(\mu)=\mu$ in the inequality (2.24), so that (2.25) still holds.